# Emergent synchrony in oscillator networks with adaptive arbitrary-order interactions

### v1

@@ Abstract
Results — As a relatively simple example of Eq. (2), a system described by equations of the form
$\frac{{\rm d}\phi_{i}}{{\rm d}t}=\omega_{i}+\frac{f_{1}(r)\epsilon_{1}}{N}\sum_{j=1}^{N}\sin(\phi_{j}-\phi_{i})+\\ \frac{f_{2}(r)\epsilon_{2}}{N^{2}}\sum_{j,k=1}^{N}\Bigg[\sin(2\phi_{k}-\phi_{j}-\phi_{i})+\xi\sin(2\phi_{k}+\phi_{j}-\phi_{i})\Bigg],$ (7)
is investigated, where $\xi=\epsilon_{2}^{\prime}/\epsilon_{2}$. For the sake of simplicity, $f_{d}(r)$ are assumed to be of the form $f_{d}(r)=(1+\gamma r)^{\pm 2}$ for $d=1$ and $d=2$, respectively, whereas $\gamma$ is a constant. In this case, Eqs. (5) and (6) take the form
$\dot{r}$ $=-\Delta r+r(1-r^{2})\Big[\sigma_{1}(r)+\sigma_{2}(r)\Big(1+\xi\cos(2\psi)\Big)\Big],$ (8)
$\dot{\psi}$ $=\omega_{0}+\xi(1+r^{2})\sigma_{2}(r)\sin(2\psi),$ (9)
respectively, where $\sigma_{n}(r)=f_{n}(r)\epsilon_{n}r^{2(n-1)}/2$ for $n=1,2$, and $\epsilon_{1}^{c}=2\Delta$. The value of $r_{s}$ can be obtained as a solution of the algebraic equation
$\left[\frac{\Delta}{1-r_{s}^{2}}-\Big(\sigma_{1}(r_{s})+\sigma_{2}(r_{s})\Big)\right]^{2}+\left[\frac{\omega_{0}}{1+r_{s}^{2}}\right]^{2}=\xi^{2}\sigma_{2}^{2}(r_{s}).$ (10)
However, in most cases, it is not possible to find closed-form analytical solutions to the above equations, thus requiring the use of numerical techniques.
For the rest of the work, $(\Delta,\xi)=(1.0,0.4)$ in Eq. (7) for the sake of simplicity. Furthermore, as an example of adaptive systems, the specific case of $\gamma=0.2$ is demonstrated in Fig. 2, for different values of $\epsilon_{1}$, whereas the value of $\omega_{0}$ and $\epsilon_{2}$ is set at $1.5$ and $5.0$, respectively  [^1] The non-adaptive case (i.e., $\gamma=0$) has already been studied in Ref. [biswas2024symmetry].. For $\epsilon_{1}=0.6$, it is evident that Eq. (10) admits no solutions. Furthermore, the asynchronous state is stable since $\epsilon_{1}<\epsilon_{1}^{c}$. Thus, the system always converges towards $r(t)\approx 0$, irrespective of initial conditions; see Fig. 2(a). However, as the value of $\epsilon_{1}$ is increased to $1.3$, there is a qualitative change in the dynamics. In this case, the asynchronous state remains stable, since $\epsilon_{1}<\epsilon_{1}^{c}$, and is evident from the temporal variation of $r(t)$, which converges towards $r(t)\approx 0$ when initialised in an asynchronous state. However, the system also demonstrates an unsteady synchronised state, where the order parameter shows oscillations about a mean value; see Fig. 2(b).
@@
Figure 3: (a)-(j): Variation of steady state order parameter $r$ as a function of $E=\epsilon_{1}/\epsilon_{2}\in[0,2]$ for different values of $\gamma$ (top-row; $\epsilon_{2}=5.0$), and $\epsilon_{2}$ (bottom-row; $\gamma=0.2$). The vertical line corresponds to the forward transition point (i.e., $E^{c}=\epsilon_{1}^{c}/\epsilon_{2}$), whereas the blue and red curves denote the forward and backward variation of $E$, respectively.
@@
As demonstrated in Fig. 2(c), increasing $\epsilon_{1}$ to $1.8$ entails further change in the dynamics. In this case, Eq. (10) admits two solutions $r_{s}=0.817$ and $r_{s}=0.849$. The eigenvalue analysis of these solutions demonstrates that the former is stable, whereas the latter is unstable in nature. However, the asynchronous state is also stable since $\epsilon_{1}<\epsilon_{1}^{c}$. Thus, the system is bistable, where the temporal variation of $r(t)$ reaches two different steady states depending on the initial condition. In particular, the asynchronous initial conditions continue to be asynchronous, whereas the synchronous initial condition converges to $r(t)\approx 0.82$. Upon further increasing $\epsilon_{1}$ to $2.6$, the asynchronous state loses stability since $\epsilon_{1}>\epsilon_{1}^{c}$, as demonstrated in Fig. 2(d). However, Eq. (10) continues to have two solutions, one at $r_{s}=0.821$, which is stable, and the other at $r_{s}=0.872$, which is unstable, as determined by the eigenvalue analysis. Thus, the system is now monostable, which is demonstrated by the temporal variation of $r(t)$, which converges around $r(t)\approx 0.82$ irrespective of the initial state of the system.
@@
Relative effect of interactions: To quantify the relative effect of pairwise and higher-order interactions, the dynamics of the system is investigated via the variation in the steady state value of the order parameter as a function of the ratio $E=\epsilon_{1}/\epsilon_{2}\in[0,2]$ in Fig. 3, representing the relative strength of the pairwise and triadic interactions. The top and bottom rows correspond to $\epsilon_{2}=5.0$ and $\gamma=0.2$, respectively. In the former, the value of $\gamma$ varies in the range $[0.05,0.25]$ in steps of $0.05$ across the sub-figures, whereas in the latter, the value of $\epsilon_{2}$ varies in the range $[2.0,6.0]$ in steps of $1.0$.
@@
In Figs. 3(a)-3(e), upon increasing $\gamma$, the transition to synchrony upon forward variation of $E$ is observed to be explosive, which is characterized by a sudden increase in $r$ at the critical point $E\approx 0.4$. In contrast, the backward transition point, captured by the sharp drop in $r$ upon reducing $E$, reduces with increasing values of $\gamma$. However, the post-transition state upon increasing $E$, as well as the route to desynchronization upon reducing $E$, varies for different values of $\gamma$. In particular, for $\gamma=0.05$, the system demonstrates an unsteady synchronous state characterized by an oscillatory temporal variation of $r(t)$ immediately after the explosive transition. Furthermore, tuning $E$ beyond $1.0$ results in a steady synchronous state; see Fig. 3(a). Subsequently, upon increasing $\gamma$ to $0.10$ and $0.15$, the region of $E$ for which the system demonstrates an unsteady synchronous state reduces, resulting in a larger parameter range for which the steady synchronous state is observed; see Figs. 3(b) and 3(c). Furthermore, upon increasing $\gamma$ to $0.20$ and $0.25$, the system exhibits only a steady synchronous state after the transition. However, a key difference is that, in the former, the system exhibits an unsteady synchronous state in a portion of the bi-stable regime before desynchronization, whereas the latter maintains the steady synchronous state enroute to desynchronization; see Figs. 3(d) and 3(e).
@@
The curves in Figs. 3(f)-3(j) demonstrate the qualitative dynamical changes upon changing $\epsilon_{2}$. Unlike the previous case, changing $\epsilon_{2}$ modifies the characteristic nature of the transition to synchrony upon varying $E$. In particular, while for $\epsilon_{2}=2.0$, the transition is observed to be continuos with a smooth increase in the steady-state value of $r$ beyond the critical point (see Fig. 3(f)), higher values of $\epsilon_{2}$ result in explosive synchronization transitions, in addition to the formation of a bi-stable zone; see Figs. 3(g)-3(j). Similar to the previous case, the post-transition synchronous states can be both unsteady or steady (for example, see Figs. 3(h) and 3(i)), whereas, for $\epsilon_{2}=6.0$, the system does not undergo any de-synchronisation transition upon reducing $E$; see Fig. 3(j).
@@
Effect of system size: Real bio-physical systems, such as the brain, are typically modelled as a connected system with a finite number of constituent dynamical units (here, oscillators), the size of which could be of different orders depending on the level of abstraction. Thus, while the previous sections demonstrate steady-state results for an infinitely large number of oscillators, the synchronization phenomena in finite-sized systems are typically affected by fluctuations which can be better characterised by constructing a probability density of the order parameter, denoted by $p(r)$. This is computed by simulating the system for a large number of time-steps and repeated over multiple initial conditions, as well as for two qualitatively different sets characterised by $r(0)\approx 0.0$ (i.e., nearly asynchronous) and $r(0)\approx 1.0$ (i.e., highly synchronous), respectively, with the results demonstrated in Fig. 4 for $N=50$.
@@
Figure 4: (a)-(d): Distribution of order parameter $p(r)$, collapsed across time, for $N=50$, and different values of $\epsilon_{1}$. The blue and red curves denote completely asynchronous and synchronous initial conditions, respectively, whereas the vertical line denotes $r_{s}$. For all figures, $(\omega_{0},\epsilon_{2},\gamma)=(1.5,5.0,0.2)$.
@@
It is observed that, for $\epsilon_{1}=0.6$ in Fig. 4(a), both distributions show a bimodal structure, with peaks near $r\approx 0.1$ and $r\approx 0.8$ respectively. However, the blue curve is substantially skewed towards $r=0$, whereas the red curve has two peaks of similar height. This indicates that, if initialized asynchronously, there is a higher probability of the system remaining asynchronous compared to transitioning to a highly synchronous state. On the other hand, if the system is initialized in a highly synchronous state, there is a nearly equal probability that the system will transition to an asynchronous state or remain in a synchronous state.
@@
In comparison, for $\epsilon_{1}=1.3$ in Fig. 4(b), both distributions retain their bimodal structure. However, in contrast to the previous case, both distributions are skewed heavily towards $r=1$. Here, the red curve is much more sharply peaked near $r\approx 0.8$ compared to $r\approx 0.1$, indicating a higher probability of finding the system in a highly synchronous state if the system is initialised in a synchronous state. In comparison, the blue curve is only slightly skewed towards the synchronous state, indicating that the probability of finding the system in a synchronous state is marginally higher than that of finding it in an asynchronous state if the system is initialised in an asynchronous state. Finally, for $\epsilon_{1}=1.8$ and $2.6$, the distributions become nearly unimodal and skewed towards $r=1$, with a peak near $r\approx 0.8$. This implies that, for any initial condition, there is a much higher probability of observing the system to be in a highly synchronous state; see Figs. 4(c) and 4(d).
@@
Conclusions — This work investigates the emergence of synchrony in a population of Kuramoto oscillators, where the interactions are both adaptive and higher-order in nature, having both rotational symmetry-preserving and symmetry-breaking components. To this end, exact analytical results for the synchronization states, their stability, and the corresponding phase transition points are obtained in the thermodynamic limit for any order of interaction. Subsequently, these were verified through large-scale simulations, focusing on the effects of triadic interactions, revealing a repertoire of dynamical states, including monostable, bistable, and unstable cases. Furthermore, varying the relative coupling strength between pairwise and higher-order interactions shows a qualitative change in the system dynamics, including a scenario where the desynchronization transition is suppressed. The effects of finite system size were also examined through numerically constructed order parameter distributions, highlighting how fluctuations in the temporal variation of the order parameter, arising out of the finite size of the system, can lead to transitions to a highly synchronous state that would not otherwise occur in the thermodynamic limit.
@@
The results presented provide important insights into the interplay of higher-order interactions and other coupling modalities, allowing for more precise delineation of transition boundaries, yielding a comprehensive “map” of the accessible states. A notable application lies in understanding epileptic seizures, where incorporating higher-order interactions could improve surgical strategies for drug-resistant epilepsy [schramm2008temporal, jirsa2023personalised, wang2023delineating] by revealing underlying epileptogenic foci that can remain obscured in purely pairwise models, thereby facilitating precise and targeted interventions. Another important example concerns the dissemination of information in social systems, where analyzing higher-order connectivity can uncover underlying “cells”, i.e., individuals or organizations that act as synchronized sources of misinformation, and could explain why misleading content often propagates far more rapidly than verified information [vosoughi2018spread]. Future work could extend this framework to involve empirical networks that include hyper-graph connectivities (for example, see Ref. [ko2022growth]), integrate domain-specific models beyond the generic Kuramoto framework adopted here, and estimate parameters using available datasets to obtain specific insights.

### v2

@@ 3 Numerical Results
In this case, Eqs. (14) and (15) take the form
$\frac{{\rm d}r}{{\rm d}t}$ $=-\Delta r+r(1-r^{2})\Big[\sigma_{1}(r)+\sigma_{2}(r)\Big(1+\xi\cos(2\psi)\Big)\Big],$ (19)
and
$\frac{{\rm d}\psi}{{\rm d}t}$ $=\omega_{0}+\xi(1+r^{2})\sigma_{2}(r)\sin(2\psi),$ (20)
respectively, where the forward phase transition point is obtained as $\epsilon_{1}^{c}=2\Delta$ from Eq. (12). Therefore, the value of $r$ in equilibrium (i.e., $r_{s}$) can be obtained as a solution of the algebraic equation
$\left[\frac{\Delta}{1-r_{s}^{2}}-\Big(\sigma_{1}(r_{s})+\sigma_{2}(r_{s})\Big)\right]^{2}+\left[\frac{\omega_{0}}{1+r_{s}^{2}}\right]^{2}=\xi^{2}\sigma_{2}^{2}(r_{s}),$ (21)
the stability of which is obtained from the Jacobian of Eqs. (19) and (20); see Eqs. (26)-(29) in A. However, finding closed-form analytical solutions is not always feasible, thus requiring the use of numerical techniques. Furthermore, even while limited to triadic interactions, the computational load for simulating Eq. (18) is extensive due to the presence of nested summations in the coupling term, especially for large system sizes. This necessitates the development of alternative schemes that reduce the computational burden, an example of which is outlined in B.
3.1 Characteristic system dynamics
For the rest of the work, $(\Delta,\xi)=(1.0,0.4)$ in Eq. (18) for the sake of simplicity. Furthermore, as an example of adaptive systems [^1] The non-adaptive case (i.e., $\gamma=0$) has already been studied in Ref. [biswas2024symmetry], facilitating direct comparison., the specific case of $\gamma=0.2$ is demonstrated in Fig. 3, for different values of $\epsilon_{1}$, whereas the value of $\omega_{0}$ and $\epsilon_{2}$ is set at $1.5$ and $5.0$, respectively. In this scenario, for $\epsilon_{1}=0.6$, it is evident that Eq. (21) admits no solutions. Furthermore, the asynchronous state is stable since $\epsilon_{1}<\epsilon_{1}^{c}$. Thus, the system always converges towards $r(t)\approx 0$, irrespective of initial conditions; see Fig. 3(a). However, as the value of $\epsilon_{1}$ is increased to $1.3$, there is a qualitative change in the dynamics. In this case, the asynchronous state remains stable, since $\epsilon_{1}<\epsilon_{1}^{c}$, and is evident from the temporal variation of $r(t)$, which converges towards $r(t)\approx 0$ when initialised in an asynchronous state. However, the system also demonstrates an unsteady synchronised state, where the order parameter shows temporal oscillations as highlighted in Fig. 3(b), a behaviour previously observed in Ref. [biswas2024symmetry] in a similar model, and also in Ref. [biswas2022mirroring] where it emerges due to the formation of partial clusters of oscillators which instantaneously overlap and separate.
@@ 3.1 Characteristic system dynamics
Figure 4: (a)-(j): Variation of the order parameter $r$ as a function of $E=\epsilon_{1}/\epsilon_{2}\in[0,2]$ for different values of $\gamma$ (top-row; $\epsilon_{2}=5.0$), and $\epsilon_{2}$ (bottom-row; $\gamma=0.2$). The vertical line corresponds to the forward transition point (i.e., $E^{c}=\epsilon_{1}^{c}/\epsilon_{2}$), whereas the blue and red curves denote the forward and backward variation of $E$, respectively, whereas $\omega_{0}=1.5$ across all figures.
@@
The observations in Fig. 3(c) demonstrate that increasing $\epsilon_{1}$ to $1.8$ entails further change in the dynamics. In this case, Eq. (21) admits two solutions $r_{s}=0.817$ and $r_{s}=0.849$. The eigenvalue analysis of these solutions demonstrates that the former is stable, whereas the latter is unstable in nature. However, the asynchronous state is also stable since $\epsilon_{1}<\epsilon_{1}^{c}$. Thus, the system is bistable, where the temporal variation of $r(t)$ reaches two different steady states depending on the initial condition. In particular, the asynchronous initial conditions continue to be asynchronous, whereas the synchronous initial condition converges to $r(t)\approx 0.82$. Finally, upon further increasing $\epsilon_{1}$ to $2.6$, the asynchronous state loses stability since $\epsilon_{1}>\epsilon_{1}^{c}$, as demonstrated in Fig. 3(d). However, Eq. (21) continues to have two solutions, one at $r_{s}=0.821$, which is stable, and the other at $r_{s}=0.872$, which is unstable, as determined by the eigenvalue analysis. Thus, the system is now monostable, which is demonstrated by the temporal variation of $r(t)$, which converges around $r(t)\approx 0.82$ irrespective of the initial state of the system.
3.2 Relative effect of interactions
@@
To quantify the relative effect of pairwise and higher-order interactions, the dynamics of the system are investigated via the variation in the steady state value of the order parameter as a function of the ratio $E=\epsilon_{1}/\epsilon_{2}\in[0,2]$ in Fig. 4, representing the relative strength of the pairwise and triadic interactions. The top and bottom rows correspond to $\epsilon_{2}=5.0$ and $\gamma=0.2$, respectively. In the former, the value of $\gamma$ varies between $[0.05,0.25]$ in steps of $0.05$ across the sub-figures, whereas in the latter, the value of $\epsilon_{2}$ varies between $[2.0,6.0]$ in steps of $1.0$.
@@
In Figs. 4(a)-4(e), upon increasing $\gamma$, the transition to synchrony upon forward variation of $E$ is observed to be explosive, which is characterised by a sudden increase in $r$ at the critical point $E\approx 0.4$. In contrast, the backward transition point, captured by the sharp drop in $r$ upon reducing $E$, reduces with increasing values of $\gamma$. However, the post-transition state upon increasing $E$, as well as the route to desynchronization upon reducing $E$, varies for different values of $\gamma$. In particular, for $\gamma=0.05$, the system demonstrates an unsteady synchronous state characterised by an oscillatory temporal variation of $r(t)$ immediately after the explosive transition. Furthermore, tuning $E$ beyond $1.0$ results in a steady synchronous state; see Fig. 4(a). Subsequently, upon increasing $\gamma$ to $0.10$ and $0.15$, the region of $E$ for which the system demonstrates an unsteady synchronous state reduces, resulting in a larger parameter range for which the steady synchronous state is observed; see Figs. 4(b) and 4(c). Furthermore, upon increasing $\gamma$ to $0.20$ and $0.25$, the system exhibits only a steady synchronous state after the transition. However, a key difference is that, in the former, the system exhibits an unsteady synchronous state in a portion of the bi-stable regime before desynchronization, whereas the latter maintains the steady synchronous state en route to desynchronization; see Figs. 4(d) and 4(e).
@@
The curves in Figs. 4(f)-4(j) demonstrate the qualitative dynamical changes upon changing $\epsilon_{2}$. Unlike the previous case, changing $\epsilon_{2}$ modifies the characteristic nature of the transition to synchrony upon varying $E$. In particular, while for $\epsilon_{2}=2.0$, the transition is observed to be continuous with a smooth increase in the steady-state value of $r$ beyond the critical point (see Fig. 4(f)), higher values of $\epsilon_{2}$ result in explosive synchronization transitions, in addition to the formation of a bi-stable zone; see Figs. 4(g)-4(j). Similar to the previous case, the post-transition synchronous states can be both unsteady or steady (for example, see Figs. 4(h) and 4(i)), whereas, for $\epsilon_{2}=6.0$, the system does not undergo any de-synchronisation transition upon reducing $E$; see Fig. 4(j).
3.3 Effect of finite system size
@@
Real bio-physical systems, such as the brain, are typically modelled as a connected system with a finite number of constituent dynamical units (here, oscillators), the size of which can vary by different orders depending on the level of abstraction. Thus, while the previous sections demonstrate steady-state results for an infinitely large number of oscillators, the synchronisation phenomena in finite-sized systems are typically affected by fluctuations which can be better characterised by constructing a probability density of the order parameter, denoted by $p(r)$. This is computed by simulating the system for a large number of time-steps (here, $10^{5}$) for a fixed set of system parameters. This is subsequently repeated over $100$ initial conditions, such that the first $50$ initial conditions correspond to $r(0)\approx 0.0$ (blue curve), and the remaining correspond to $r(0)\approx 1.0$ (red curve), with the results for $N=50$ shown in Fig. 5.
@@
Figure 5: (a)-(d): Numerically constructed distribution of the order parameter, i.e., $p(r)$, collapsed across time, for $N=50$, and different values of $\epsilon_{1}$. The blue and red curves denote completely asynchronous and synchronous initial conditions, respectively, whereas the vertical line denotes $r_{s}$. For all figures, $(\omega_{0},\epsilon_{2},\gamma)=(1.5,5.0,0.2)$.
@@
It is observed that, for $\epsilon_{1}=0.6$ in Fig. 5(a), both distributions show a bimodal structure, with peaks near $r\approx 0.1$ and $r\approx 0.8$, respectively. However, the blue curve is substantially skewed towards $r=0$, whereas the red curve has two peaks of similar height. This indicates that, if initialised asynchronously, there is a higher probability of the system remaining asynchronous compared to transitioning to a highly synchronous state. On the other hand, if the system is initialised in a highly synchronous state, there is a nearly equal probability that the system will transition to an asynchronous state or remain in a highly synchronous state.
@@
In comparison, for $\epsilon_{1}=1.3$ in Fig. 5(b), both distributions retain their bimodal structure. However, in contrast to the previous case, both distributions are skewed heavily towards $r=1$. Here, the red curve is much more sharply peaked near $r\approx 0.8$ compared to $r\approx 0.1$, indicating a higher probability of finding the system in a highly synchronous state if the system is initialised in a synchronous state. In comparison, the blue curve is only slightly skewed towards the synchronous state, indicating that the probability of finding the system in a synchronous state is marginally higher than that of finding it in an asynchronous state if the system is initialised in an asynchronous state. Finally, for $\epsilon_{1}=1.8$ and $2.6$, the distributions become nearly unimodal and skewed towards $r=1$, with a peak near $r\approx 0.8$. This implies that, for any initial condition, there is a much higher probability of observing the system to be in a highly synchronous state; see Figs. 5(c) and 5(d). Taken together, the results presented in Fig. 5 highlight the effects of size-induced fluctuations in driving the system to emergent states not observed in the infinite-size limit. Furthermore, additional plots of $p(r)$, obtained by repeating the same procedure for different values of $\epsilon_{1}$ and $N$, are presented in Fig. 6 of C, which highlight the qualitative variation in the characteristic order parameter dynamics as a function of system size.
4 Conclusions & Future Directions
@@
This work presents a very general theory of synchronization in a population of Kuramoto oscillators, where the interactions have both rotational symmetry-preserving and symmetry-breaking components. Furthermore, the additional terms are included such that they capture a wide range of coupling mechanisms, introduced through adaptation functions that modify the instantaneous coupling strength as a function of the synchronization order parameter, as well as the effects of frustration incorporated using a constant phase lag. Although these factors have been studied individually in the literature, this work introduces a unified mathematical framework to explore their collective interplay. To this end, exact analytical results for the synchronisation states, their stability, and the corresponding phase transition points are obtained in the thermodynamic limit for any order of interaction using the popular Ott-Antonsen Ansatz. Subsequently, these analytical results were verified through large-scale numerical simulations, focusing on the combined effects of pairwise and triadic interactions and an adaptation function having a quadratic dependence on the order parameter, revealing a repertoire of dynamical states, including monostable, bistable, and unstable synchronous cases. Furthermore, varying the relative coupling strength between pairwise and higher-order interactions shows a qualitative change in the system dynamics, highlighting both continuous and explosive phase transitions, as well as a scenario where the desynchronization transition is suppressed. Finally, the effects of finite system size were also examined through numerically constructed order parameter distributions. The obtained results highlight how fluctuations in the temporal variation of the order parameter, arising out of the finite size of the system, can lead to transitions to a highly synchronous state that would not otherwise occur in the thermodynamic limit. Such transitions occur spontaneously, i.e., without any change in system parameters, and are distinct from those observed in a system of nearly-infinite size, where transitions occur only upon tuning system parameters across a critical transition point.
@@
The results presented provide important insights into the interplay of higher-order interactions and other coupling modalities, allowing for more precise delineation of transition boundaries, yielding a comprehensive “map” of the accessible states. Such models find widespread application across domains, among which a notable application lies in understanding epileptic seizures, where incorporating higher-order interactions of different orders could improve surgical strategies for drug-resistant epilepsy [schramm2008temporal, jirsa2023personalised, wang2023delineating] by revealing underlying epileptogenic foci that can remain obscured in purely pairwise models, thereby facilitating precise and targeted interventions. Another important example concerns the dissemination of information in social systems, where analysing higher-order connectivity can uncover underlying “cells”, i.e., individuals or organisations that act as synchronised sources of misinformation, and could explain why misleading content often propagates far more rapidly than verified information [vosoughi2018spread].
However, in spite of the theoretical advancements presented in this study, certain key limitations still remain. This includes obtaining the exact order of interaction driving the explosive spread of synchronous activity in various physical systems, as well as the explicit functional form of the adaptation functions governing the overall order parameter dynamics, thereby opening up many potential directions of future research. Furthermore, this will also include extending this framework to involve empirical hyper-graph connectivities (for example, see Ref. [ko2022growth]), integrating domain-specific models beyond the generic Kuramoto model, and estimating the various systems parameters using available datasets to obtain problem-specific insights.
Author Contributions
DB: Conceptualization (equal); Formal analysis (equal); Investigation (lead); Methodology (equal); Software (lead); Validation (equal); Visualization (lead); Writing – original draft (equal). AB: Conceptualization (equal); Funding acquisition (lead); Investigation (supporting); Methodology (supporting); Project administration (lead); Supervision (lead); Visualization (supporting); Writing – review & editing (lead).
Acknowledgements